Prev Next

SW&IT Meeting Minutes 2020-11-23

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20201116 - Bec, Matthieu 40.0
LR20201116 - Holmberg, Carl 40.0
LR20201116 - Edmeades, David 0.0
LR20201116 - Sola, Igor 0.0
LR20201116 - Grenz, Paul 4.0
LR20201116 - Hooper, Stephen 80.0
LR20201116 - Bryant, Trey 25.0
LR20201116 - Zhang, Yang 40.0
LR20201116 - Zhang, Xianyu 40.0

News from management

  • staff announcements
  • Thu26/Fri27 Nov Holidays

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports


  • Last week
    • OC, Management
    • ATT SW, ATT PM monthly, ROWG, SHARK-VIS waffle, AO MCS SW meetings
    • DMS AO discussion with SOUL team
    • cobbler demo to IT
    • iLocater fiber discussions
  • This week
    • ATT SW
    • PP2020


  • Last week:
    • AGW:
      • Delta Tau tech support responds that their PMAC API does not expose firmware download to the user, and we’re stuck with PEWIN32 for this purpose. Intern summer project to reverse engineer the undocumented API?
      • Pushed new firmware makefiles and configuration modules to oacontrol, supporting combos of old/new PMACs, with grey/binary encoders and interface boards.
      • JP finds that the firmware code we thought calculated the encoder angle for all movements is actually for initialization during the OAC ‘home’ command. We need to document the timeline of read/writes from the command ‘home -m21’ to sample the correct moment the angle is calculated.
    • AHS: Added initialization parameter status check to LBTI status script.
    • statserv: Eliminated all INDI drivers except indialh and ovms from indiserver startup arguments. Others were superfluous.
  • This week:
    • AGW: Assist JP in creating timeline when firmware Q40, Q41, and Q150 registers are modified by oac commands (ex. ‘Home’).
    • AHS: Add collection of _STATE parameter to validate reading Cryomech ADAM data from LBTI INDI service.
    • statserv: Add LBTI and ARGOS overall status updates.


  • UPS batteries ordered; will arrive today
    • Will plan office trip to work on battery modules, likely next week.
  • Couple of phishing alerts from Google; followed up with user to verify that he deleted them
  • Remind your families about phishing and advance fee scams. The holiday push is upon us.
  • Archive backup got stuck, reboot fixed
  • Given resource analysis on archive-in, we decided to deallocate CPU and RAM; everything seems to be working.
  • UA IT training: don't give out your password!!!!
  • IBM storage presentation
  • AB laptop likely to be released, pending discussion with him and Sciops.


  • discovered the long suspected flag to make sed overwrite files while using regex to substitute text exists - *-i* flag
  • did a bunch of updates to M1M3 SS, ask Carl or Yang for details (better logging, QPlainTextEdit for log view,..)


  • Last week
    • Wrote design document for white light source for the ARGOS swing arm.
    • Forked tcs code to begin working on implementing the new code.
  • This week
    • Implementing the code for the swing arm and white light source.


  • Last week
    • ASM elevation limit display changes on AO supervisor side
    • AOS/ALH for TSS display invistigation
    • ATT AdSec
    • ATT SW
    • AO operation support
    • Anemometer disconnection (voltage unstable, communication loss) investigation
  • This week


  • Previous:
    • Usual TMS, Cravat/OSCO work and meeting.
    • Some misc work (MODS data cleaning script etc).
  • This week:
    • Finish some TMS work (SPIE paper section, some requested documents and software changes).
    • modsAlign refactoring.


  • DIMM hardware
    • new hardware quote
  • measurement of mountain bandwidth
  • scripting for SSL Let's Encrypt
  • MAC address registration * UITS configuration (incomplete on their side)
  • LBC hardwired paths
  • AllSky mixed SSL
  • Laptop discussion, IT meeting
  • ROWG monitors
  • DIMM hardware
  • PEPSI, obs, VM with Leroy
  • LBC cart


Last week:
  • LUCI
    • RPMs generation.
    • Geirs script followed up on Barry's night log.
  • PIT
    • I looked at the issue reported by a user where a proposal was submitted and we got no record of it. Given the implication of it, the fact that I could not find a way for the code to behave in this way I wrote a report for others to chip in if they'd like and see if they agree or not with my assessment. PIT submit issue report

This week:
  • Harassment training
  • LUCI
    • RPM
    • IT 8231 GEIRS timeout
    • IT 8220
    • split current luci motor privilege into two to allow more control on what people can and can't do.
    • Look into reading WebIO status and passing it to the database/telemetry.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
  • PIT
    • pit admin tool
Topic revision: r10 - 23 Nov 2020, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback